# Non-reversible lifts of reversible diffusion processes and relaxation times

### 07 Feb 2024 v1

@@ 2 Lifts of reversible diffusions
In the following, we consider a reversible diffusion process with state space $\mathcal{S}$ and a time-homogenous Markov process with state space $\hat{\mathcal{S}}=\mathcal{S}\times\mathcal{V}$ and invariant probability measures $\mu(\mathop{}\!\mathrm{d}x)$ and $\hat{\mu}(\mathop{}\!\mathrm{d}x\mathop{}\!\mathrm{d}v)=\mu(\mathop{}\!\mathrm{d}x)\kappa(x,\mathop{}\!\mathrm{d}v)$, respectively. We frequently think of the components of the latter as a position and a velocity or momentum variable. The associated Markov semigroups acting on $L^{2}(\mu)$ and $L^{2}(\hat{\mu})$ are $(P_{t})_{t\geq 0}$ and $(\hat{P}_{t})_{t\geq 0}$, and their infinitesimal generators are denoted by $(L,\operatorname{Dom}(L))$ and $(\hat{L},\operatorname{Dom}(\hat{L}))$, respectively. While the underlying diffusion is assumed to be reversible, the second process will not be reversible in general, as an improvement in convergence rate is usually achieved through non-reversibility.
@@
$\operatorname{Dom}(L)\circ\pi=\{g\circ\pi\colon g\in\operatorname{Dom}(L)\}\subseteq\operatorname{Dom}(\hat{L})$ (6)
@@
$\frac{1}{2}\int_{\hat{\mathcal{S}}}\hat{L}(f\circ\pi)\hat{L}(g\circ\pi)\mathop{}\!\mathrm{d}\hat{\mu}=-\int_{\mathcal{S}}fLg\mathop{}\!\mathrm{d}\mu\,.$ (8)
We also say that $(\hat{L},\operatorname{Dom}(\hat{L}))$ is a second-order lift of $(L,\operatorname{Dom}(L))$. Since we only consider second-order lifts, we may also simply refer to second-order lifts as lifts.
@@
    In 1, it is sufficient to verify the two conditions for functions $f$ and $g$ from a core of the generator $(L,\operatorname{Dom}(L))$ or even from a core of the associated Dirichlet form
    $\mathcal{E}(f,g)=-\int_{\mathcal{S}}fLg\mathop{}\!\mathrm{d}\mu\,.$
    Similarly, it is sufficient to check that for all $f\in\operatorname{Dom}(L)$ and $x\in\mathcal{S}$
@@
    $\frac{1}{2}\int_{\hat{\mathcal{S}}}\left(\hat{L}(f\circ\pi)\right)^{2}\mathop{}\!\mathrm{d}\hat{\mu}=-\int_{\mathcal{S}}fLf\mathop{}\!\mathrm{d}\mu$
@@
    $\int_{\hat{\mathcal{S}}}(f\circ\pi)\hat{L}(g\circ\pi)\mathop{}\!\mathrm{d}\hat{\mu}=\int_{\mathcal{S}}fLg\mathop{}\!\mathrm{d}\mu$
@@ 4 Lower bounds for relaxation times of lifts
Let $\lambda>\operatorname{gap}(L)$. By the variational characterisation of the spectral gap, there exists $f\in\operatorname{Dom}(L)\cap L_{0}^{2}(\mu)$ such that $(f,-Lf)\leq\lambda\lVert f\rVert_{L^{2}(\mu)}^{2}$. By 1, $f\circ\pi\in\operatorname{Dom}(\hat{L})\cap L_{0}^{2}(\hat{\mu})$ and
$\lVert\hat{L}(f\circ\pi)\rVert_{L^{2}(\hat{\mu})}^{2}=2(f,-Lf)\leq 2\lambda\lVert f\rVert_{L^{2}(\mu)}^{2}=2\lambda\lVert f\circ\pi\rVert_{L^{2}(\hat{\mu})}^{2}\,.$
By considering the limit $\lambda\to\operatorname{gap}(L)$, we see that $s(\hat{L})\leq\sqrt{2\operatorname{gap}(L)}$, which implies the assertion by Lemma 10. ∎
@@
It is an open problem whether also $\frac{1}{\nu}$ can be lower bounded in terms of $\frac{1}{\sqrt{\operatorname{gap}(L)}}$. This would follow from an upper bound on $\operatorname{gap}(\hat{L})$ by a multiple of $\sqrt{\operatorname{gap}(L)}$. We have not managed to prove such a bound. This is similar to the discrete time case, where a lower bound for the mixing time is available but an upper bound for the spectral gap of lifts is not known [15].
@@ 5.1 Space-time Poincaré inequalities and contractivity
where $\lambda|_{[0,T]}$ is the Lebesgue measure on $[0,T]$. Furthermore, we consider the operator $(A,\operatorname{Dom}(A))$ on $L^{2}(\overline{\mu}\otimes\kappa)$ defined by
@@
$\lVert f-\int f\mathop{}\!\mathrm{d}(\overline{\mu}\otimes\kappa)\rVert_{L^{2}(\overline{\mu}\otimes\kappa)}^{2}\leq C_{0}(T)\lVert f-\Pi_{v}f\rVert_{L^{2}(\overline{\mu}\otimes\kappa)}^{2}+C_{1}(T)\lVert Af\rVert_{L^{2}(\overline{\mu}\otimes\kappa)}^{2}\,.$
@@
$\nu=\frac{\gamma}{C_{0}(T)+\gamma^{2}C_{1}(T)}\,.$
The choice $\gamma=\sqrt{\frac{C_{0}(T)}{C_{1}(T)}}$ maximises the rate and gives $\nu=\frac{1}{2\sqrt{C_{0}(T)C_{1}(T)}}$. For this choice,
@@
$\lVert\hat{P}_{t+\cdot}^{(\gamma)}f\rVert_{L^{2}(\overline{\mu}\otimes\kappa)}^{2}\leq(C_{0}(T)+\gamma^{2}C_{1}(T))\lVert(\Pi_{v}-I)\hat{P}_{t+\cdot}^{(\gamma)}f\rVert_{L^{2}(\overline{\mu}\otimes\kappa)}^{2}\,,$
@@
$\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}t}\int_{t}^{t+T}\lVert\hat{P}_{s}^{(\gamma)}f\rVert_{L^{2}(\hat{\mu})}^{2}\mathop{}\!\mathrm{d}s\leq\frac{-2\gamma}{C_{0}(T)+\gamma^{2}C_{1}(T)}\int_{t}^{t+T}\lVert\hat{P}_{s}^{(\gamma)}f\rVert_{L^{2}(\hat{\mu})}^{2}\mathop{}\!\mathrm{d}s\,.$
@@
$\frac{1}{T}\int_{t}^{t+T}\lVert\hat{P}_{s}^{(\gamma)}f\rVert_{L^{2}(\hat{\mu})}^{2}\mathop{}\!\mathrm{d}s$ $\leq\exp\left(\frac{-2\gamma}{C_{0}(T)+\gamma^{2}C_{1}(T)}t\right)\frac{1}{T}\int_{0}^{T}\lVert\hat{P}_{s}^{(\gamma)}f\rVert_{L^{2}(\hat{\mu})}^{2}\mathop{}\!\mathrm{d}s$
@@ 5.4 Space-time Poincaré inequalities via lifts
We consider the adjoint $(A^{*},\operatorname{Dom}(A^{*}))$ in $L^{2}([0,T]\times\mathbb{R}^{d},\overline{\mu}\otimes\kappa)$ of the operator $(A,\operatorname{Dom}(A))$ defined in (22). By integration by parts and the antisymmetry of $\hat{L}$,
@@
The following proof is an adapted, optimised, and partially simplified version of the proof of [36, Theorem 2.1].
@@ 5.5 Optimality of randomised Hamiltonian Monte Carlo
$\nu=\frac{\gamma}{2c_{0}(T)+(3+4c_{1}(T))\gamma^{2}}\,,$ (35)
@@
For fixed $T\in(0,\infty)$, the rate $\nu$ in (35) is maximal when $\gamma$ is chosen such that the two summands in the denominator coincide, i.e. for $\gamma=\sqrt{2c_{0}(T)/(3+4c_{1}(T))}$. In this case,
@@
$\frac{1}{\nu}=\frac{482}{\sqrt{m}}\left(\frac{1}{\sqrt{m}\gamma}+\left(2124\tfrac{1}{3}+300\frac{\kappa_{-}}{m}\right)\frac{\sqrt{m}\gamma}{482}\right)<\frac{482}{\sqrt{m}}\left(\frac{1}{\sqrt{m}\gamma}+5\left(1+\frac{\kappa_{-}}{7m}\right){\sqrt{m}\gamma}\right)\,.$ (37)
@@
$\gamma=\sqrt{482/(2124\tfrac{1}{3}\,m+300\,\kappa_{-})}\,,$ (38)
@@
$\frac{1}{\nu}=2\cdot\frac{482}{m\gamma}<\frac{2024}{\sqrt{m}}\cdot\sqrt{1+\frac{\kappa_{-}}{7m}}\,.$ (39)
@@
for any $\gamma$ satisfying $1/A\leq\sqrt{m}\gamma\leq A$.

### 07 Mar 2024 v3

@@ 2 Lifts of reversible diffusions
In the following, we consider a reversible diffusion process with state space $\mathcal{S}$ and a time-homogenous Markov process with state space $\hat{\mathcal{S}}=\mathcal{S}\times\mathcal{V}$ and invariant probability measures $\mu(\mathop{}\!\mathrm{d}x)$ and $\hat{\mu}(\mathop{}\!\mathrm{d}x\mathop{}\!\mathrm{d}v)=\mu(\mathop{}\!\mathrm{d}x)\kappa(x,\mathop{}\!\mathrm{d}v)$, respectively. We frequently think of the components of the latter as a position and a velocity or momentum variable. The associated Markov semigroups acting on $L^{2}(\mu)$ and $L^{2}(\hat{\mu})$ are $(P_{t})_{t\geq 0}$ and $(\hat{P}_{t})_{t\geq 0}$, and their infinitesimal generators are denoted by $(L,\operatorname{Dom}(L))$ and $(\hat{L},\operatorname{Dom}(\hat{L}))$, respectively. While the underlying diffusion is assumed to be reversible, the second process will not be reversible in general, as an improvement in convergence rate is usually achieved through non-reversibility. The Dirichlet form associated to $(L,\operatorname{Dom}(L))$ is the extension of
$\mathcal{E}(f,g)=-\int_{\mathcal{S}}fLg\mathop{}\!\mathrm{d}\mu.$
to a closed symmetric bilinear form with domain $\operatorname{Dom}(\mathcal{E})$ given by the closure of $\operatorname{Dom}(L)$ with respect to the norm $\lVert f\rVert_{L^{2}(\mu)}+\mathcal{E}(f,f)^{1/2}$.
@@
    $\operatorname{Dom}(L)\subseteq\{f\in L^{2}(\mu)\colon f\circ\pi\in\operatorname{Dom}(\hat{L})\}$ (6)
@@
    $\frac{1}{2}\int_{\hat{\mathcal{S}}}\hat{L}(f\circ\pi)\hat{L}(g\circ\pi)\mathop{}\!\mathrm{d}\hat{\mu}=\mathcal{E}(f,g)\,.$ (8)
  2. (ii)
    The semigroup $(\hat{P}_{t})_{t\geq 0}$ is a strong second-order lift of $(P_{t})_{t\geq 0}$ if
    $\operatorname{Dom}(\mathcal{E})=\{f\in L^{2}(\mu)\colon f\circ\pi\in\operatorname{Dom}(\hat{L})\}$ (9)
    and (7) and (8) are satisfied for all $f,g\in\operatorname{Dom}(\mathcal{E})$.
We also say that $(\hat{L},\operatorname{Dom}(\hat{L}))$ is a (strong) second-order lift of $(L,\operatorname{Dom}(L))$. Since we only consider second-order lifts, we may also simply refer to second-order lifts as lifts. In Theorem 11 below, we will give a lower bound of the relaxation time that holds for arbitrary second-order lifts.
@@
    In general, a non-reversible Markov process can be a lift in the sense of 1 (i) of several reversible diffusion processes (for example with different boundary behaviours). On the other hand, if $(\hat{P}_{t})_{t\geq 0}$ is a strong lift of $(P_{t})_{t\geq 0}$, then the Dirichlet form $(\mathcal{E},\operatorname{Dom}(\mathcal{E}))$ and, correspondingly, the semigroup $(P_{t})_{t\geq 0}$ are uniquely determined by $(\hat{P}_{t})_{t\geq 0}$ through the conditions (8) and (9). In this case, the semigroup $(P_{t})_{t\geq 0}$ is called the collapse of $(\hat{P}_{t})_{t\geq 0}$.
  2. (ii)
    In 1, it is sufficient to verify conditions (7) and (8) for functions $f$ and $g$ from a core of the generator $(L,\operatorname{Dom}(L))$ or from a core of the associated Dirichlet form. Similarly, it is sufficient to check that
@@
    $\frac{1}{2}\int_{\hat{\mathcal{S}}}\left(\hat{L}(f\circ\pi)\right)^{2}\mathop{}\!\mathrm{d}\hat{\mu}=\mathcal{E}(f,f)$
@@
    $\int_{\hat{\mathcal{S}}}(f\circ\pi)\hat{L}(g\circ\pi)\mathop{}\!\mathrm{d}\hat{\mu}=\int_{\mathcal{S}}fLg\mathop{}\!\mathrm{d}\mu$ (10)
@@ 4 Lower bounds for relaxation times of lifts
Let $\lambda>\operatorname{gap}(L)$. By the variational characterisation of the spectral gap, there exists $f\in\operatorname{Dom}(L)\cap L_{0}^{2}(\mu)$ such that $(f,-Lf)_{L^{2}(\mu)}\leq\lambda\lVert f\rVert_{L^{2}(\mu)}^{2}$. By 1, $f\circ\pi\in\operatorname{Dom}(\hat{L})\cap L_{0}^{2}(\hat{\mu})$ and
$\lVert\hat{L}(f\circ\pi)\rVert_{L^{2}(\hat{\mu})}^{2}=2(f,-Lf)_{L^{2}(\mu)}\leq 2\lambda\lVert f\rVert_{L^{2}(\mu)}^{2}=2\lambda\lVert f\circ\pi\rVert_{L^{2}(\hat{\mu})}^{2}\,.$ (22)
By considering the limit $\lambda\to\operatorname{gap}(L)$, we see that $\operatorname{s}(\hat{L})\leq\sqrt{2\operatorname{gap}(L)}$, which implies the assertion by Lemma 10. ∎
@@
  1. (i)
    Consider a Markov chain in discrete or continuous time with state space $\hat{\mathcal{S}}$, transition kernel $\hat{p}$ and invariant measure $\hat{\mu}$ that is a first-order lift (in the sense of [15], or equivalently, (10)) of a reversible Markov chain with state space $\mathcal{S}$, transition kernel $p$ and invariant measure $\mu$. That is,
    $(f\circ\pi,\hat{p}(g\circ\pi))_{L^{2}(\hat{\mu})}=(f,pg)_{L^{2}(\mu)}$
    for all bounded measurable functions $f,g\colon\mathcal{S}\to\mathbb{R}$. Then the associated generators $\hat{L}=\hat{p}-I$ and $L=p-I$ satisfy
    $\lVert\hat{L}(f\circ\pi)\rVert_{L^{2}(\hat{\mu})}^{2}$ $=\lVert(\hat{p}-I)(f\circ\pi)\rVert_{L^{2}(\hat{\mu})}^{2}$
    $=\lVert f\circ\pi\rVert_{L^{2}(\hat{\mu})}^{2}-2(f\circ\pi,\hat{p}(f\circ\pi))_{L^{2}(\hat{\mu})}+\lVert\hat{p}(f\circ\pi)\rVert_{L^{2}(\hat{\mu})}^{2}$
    $\leq 2\lVert f\rVert_{L^{2}(\mu)}^{2}-2(f,pf)_{L^{2}(\mu)}=2(f,-Lf)_{L^{2}(\mu)}\,,$
    since $\hat{p}$ is a contraction. Hence the second-order lift condition (8) holds with an inequality, so that we can conclude $\operatorname{s}(\hat{L})\leq\sqrt{2\operatorname{gap}(L)}$ similarly as in (22). Lemma 10 can then be used to obtain a lower bound on the relaxation time of the lifted Markov chain in terms of the spectral gap of $L$ as in Theorem 11.
  2. (ii)
    It is an open problem whether also $\frac{1}{\nu}$ can be lower bounded in terms of $1/\sqrt{\operatorname{gap}(L)}$. This would follow from an upper bound on $\operatorname{gap}(\hat{L})$ by a multiple of $\sqrt{\operatorname{gap}(L)}$. We have not managed to prove such a bound. This is similar to the discrete time case, where a lower bound for the mixing time is available but an upper bound for the spectral gap of lifts is not known [15].
@@ 5.1 Space-time Poincaré inequalities and contractivity
where $\lambda|_{[0,T]}$ is the Lebesgue measure on $[0,T]$. Furthermore, we consider the operator $(A,\operatorname{Dom}(A))$ on $L^{2}([0,T]\times\mathbb{R}^{d}\times\mathbb{R}^{d},\overline{\mu}\otimes\kappa)$ defined by
@@
$\lVert f-\int f\mathop{}\!\mathrm{d}(\overline{\mu}\otimes\kappa)\rVert_{L^{2}(\overline{\mu}\otimes\kappa)}^{2}\leq C_{0}(T)\lVert Af\rVert_{L^{2}(\overline{\mu}\otimes\kappa)}^{2}+C_{1}(T)\lVert f-\Pi_{v}f\rVert_{L^{2}(\overline{\mu}\otimes\kappa)}^{2}\,.$
@@
$\nu=\frac{\gamma}{\gamma^{2}C_{0}(T)+C_{1}(T)}\,.$
The choice $\gamma=\sqrt{\frac{C_{1}(T)}{C_{0}(T)}}$ maximises the rate and gives $\nu=\frac{1}{2\sqrt{C_{0}(T)C_{1}(T)}}$. For this choice,
@@
$\lVert\hat{P}_{t+\cdot}^{(\gamma)}f\rVert_{L^{2}(\overline{\mu}\otimes\kappa)}^{2}\leq(\gamma^{2}C_{0}(T)+C_{1}(T))\lVert(\Pi_{v}-I)\hat{P}_{t+\cdot}^{(\gamma)}f\rVert_{L^{2}(\overline{\mu}\otimes\kappa)}^{2}\,,$
@@
$\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}t}\int_{t}^{t+T}\lVert\hat{P}_{s}^{(\gamma)}f\rVert_{L^{2}(\hat{\mu})}^{2}\mathop{}\!\mathrm{d}s\leq\frac{-2\gamma}{\gamma^{2}C_{0}(T)+C_{1}(T)}\int_{t}^{t+T}\lVert\hat{P}_{s}^{(\gamma)}f\rVert_{L^{2}(\hat{\mu})}^{2}\mathop{}\!\mathrm{d}s\,.$
@@
$\frac{1}{T}\int_{t}^{t+T}\lVert\hat{P}_{s}^{(\gamma)}f\rVert_{L^{2}(\hat{\mu})}^{2}\mathop{}\!\mathrm{d}s$ $\leq\exp\left(\frac{-2\gamma}{\gamma^{2}C_{0}(T)+C_{1}(T)}t\right)\frac{1}{T}\int_{0}^{T}\lVert\hat{P}_{s}^{(\gamma)}f\rVert_{L^{2}(\hat{\mu})}^{2}\mathop{}\!\mathrm{d}s$
@@ 5.4 Space-time Poincaré inequalities via lifts
We consider the adjoint $(A^{*},\operatorname{Dom}(A^{*}))$ in $L^{2}([0,T]\times\mathbb{R}^{d}\times\mathbb{R}^{d},\overline{\mu}\otimes\kappa)$ of the operator $(A,\operatorname{Dom}(A))$ defined in (25). By integration by parts and the antisymmetry of $\hat{L}$,
@@
The following proof is an adapted, optimised, and partially simplified version of the proof of [38, Theorem 2.1].
@@ 5.5 Optimality of randomised Hamiltonian Monte Carlo
$\nu=\frac{\gamma}{2c_{0}(T)\gamma^{2}+3+4c_{1}(T)}\,,$ (38)
@@
For fixed $T\in(0,\infty)$, the rate $\nu$ in (38) is maximal when $\gamma$ is chosen such that the two summands in the denominator coincide, i.e. for $\gamma=\sqrt{(3+4c_{1}(T))/(2c_{0}(T))}$. In this case,
@@
$\frac{1}{\nu}=\frac{482}{\sqrt{m}}\left(\frac{\gamma}{\sqrt{m}}+\left(2124\tfrac{1}{3}+300\frac{\kappa_{-}}{m}\right)\frac{\sqrt{m}}{482\gamma}\right)<\frac{482}{\sqrt{m}}\left(\frac{\gamma}{\sqrt{m}}+5\left(1+\frac{\kappa_{-}}{7m}\right){\frac{\sqrt{m}}{\gamma}}\right)\,.$ (40)
@@
$\gamma=\sqrt{(2124\tfrac{1}{3}\,m+300\,\kappa_{-})/482}\,,$ (41)
@@
$\frac{1}{\nu}=2\cdot\frac{482\gamma}{m}<\frac{2024}{\sqrt{m}}\cdot\sqrt{1+\frac{\kappa_{-}}{7m}}\,.$ (42)
@@
for any $\gamma$ satisfying $1/A\leq\frac{\gamma}{\sqrt{m}}\leq A$.
@@
Remark 29 (Hypocoercivity approaches).
Based on the work of Hérau [29], Dolbeault, Mouhot and Schmeiser [19, 20] developed a method for proving hypocoercivity for a general class of linear kinetic equations, covering the kinetic Fokker-Planck and the linear Boltzmann equations and thereby Langevin dynamics and randomised HMC. Their method is based on a modification of the $L^{2}(\mu\otimes\kappa)$-norm by a suitable perturbation which allows to recover dissipation in all variables. As pointed out to us by Lihan Wang, if $\hat{L}$ is the generator of the Hamiltonian flow, then Assumption H3 in [20] resembles our first-order condition (7) for a second-order lift, while Assumption H2 in [20] is implied by the second-order condition (8) together with 2(ii) on the spectral gap of the underlying diffusion. Since the entropy functional in [20] does not directly capture the dynamics of the process, one cannot expect optimal bounds on the convergence rate. Indeed, as calculated in [13, Appendix B], assuming a lower bound on the Hessian of the potential and in the case that $\gamma$ is proportional to $\sqrt{m}$, the approach in [20] seems to yield a rate of order $m^{5/2}$ in the limit as $m\to 0$ instead of the optimal order $m^{1/2}$.
Statements and Declarations
Funding.    Gefördert durch die Deutsche Forschungsgemeinschaft (DFG) im Rahmen der Exzellenzstrategie des Bundes und der Länder – GZ2047/1, Projekt-ID 390685813.
The authors were funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – GZ 2047/1, Project-ID 390685813.
Acknowledgements.    The authors would like to thank Lihan Wang, Arnaud Guillin and Léo Hahn for many helpful discussions.